# SPIRIT: Short-term Prediction of solar IRradIance for zero-shot Transfer learning using Foundation Models

### v1

@@ Abstract.
Traditional solar forecasting models are based on several years of site-specific historical irradiance data, often spanning five or more years, which are unavailable for newer photovoltaic farms. As renewable energy is highly intermittent, building accurate solar irradiance forecasting systems is essential for efficient grid management and enabling the ongoing proliferation of solar energy, which is crucial to achieve the United Nations’ net zero goals. In this work, we propose SPIRIT, a novel approach leveraging foundation models for solar irradiance forecasting, making it applicable to newer solar installations. Our approach outperforms state-of-the-art models in zero-shot transfer learning by about 70%, enabling effective performance at new locations without relying on any historical data. Further improvements in performance are achieved through fine-tuning, as more location-specific data becomes available. These findings are supported by statistical significance, further validating our approach. SPIRIT represents a pivotal step towards rapid, scalable, and adaptable solar forecasting solutions, advancing the integration of renewable energy into global power systems.
Solar Forecasting, Renewable Energy, Foundation Models, Transfer Learning, Zero-shot Learning, Fine-tuning, Deep Learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; 978-1-4503-XXXX-X/18/06isbn: ;ccs: Computing methodologies Machine learningccs: Applied computing Forecastingccs: Hardware Renewable energyccs: Computing methodologies Foundation modelsccs: Computing methodologies Transfer learning
@@
Figure 1. Illustration of our system: A vision encoder (top-left) extracts embeddings from a sky camera image sampled from a diverse set spanning multiple locations and setups. Physics-inspired features are derived and integrated with auxiliary values, then merged with the image embedding (top-middle) into a unified representation. For nowcasting (right), a regressor predicts Global Horizontal Irradiance from this feature vector. For forecasting (bottom), a time-series model processes past feature vectors to create a context embedding, which is concatenated with a future covariate vector—constructed from known future values—to form the final latent representation. A regressor then maps this representation to future GHI values (bottom-right).
1. Introduction
@@
The proliferation of solar energy is paramount for electrification and the global energy transition to meet the Net Zero commitments of the United Nations (Sadhukhan, 2022). As the world moves toward renewable sources, solar energy is notable for its accessibility and potential to significantly reduce carbon emissions (Sen, 2008). Expanding the solar energy infrastructure is crucial to mitigate the effects of climate change (Bashir et al., 2021) and meet the energy demands arising from sectors such as data centers (Agarwal et al., 2021), transportation (Lee et al., 2016), and buildings (Iyengar et al., 2017).
@@
Unlike conventional power sources such as thermal and nuclear, solar energy has inherent shortcomings. Its intermittency, due to the daily and seasonal variations in sunlight, poses significant challenges for energy grid stability (Abido et al., 2022). One notable issue arising from the higher penetration of solar power is the “duck curve” (Iyengar et al., 2016), where the mismatch between solar energy production and peak energy demand leads to significant challenges in grid management. Although storage capacity is increasing, electricity grids typically operate as a just-in-time system where energy supply and demand must be balanced (Joskow, 2012). To ensure grid efficiency, renewable operators must pay a deviation penalty to discourage unplanned energy contributions, thereby maintaining a balanced and predictable energy supply (Yang et al., 2020). Thus, accurate short-term solar predictions are crucial for the efficient operation of the energy grid (Iyengar et al., 2014).
@@
Existing approaches for short-term forecasting use Sky Cameras — i.e., a fish-eye lens camera positioned to look directly towards the zenith — which require extensive site-specific data to train models (Hammond et al., 2024; Gao and Liu, 2022). These models have demonstrated the ability to develop high-accuracy models, albeit using training data spanning multiple years. With the overall solar PV fleet expected to increase from 1 TW in 2022 to 10 TW by 2030 (isa, 2023), 90% of the solar farms worldwide will have negligible data to train custom models from scratch. Thus, lack of sufficient site-specific solar data underscores the need for approaches that do not compromise model performance.
@@
With the advent of vision foundation models, we have seen improvement in accuracy of various Computer Vision tasks — such as feature extraction, object detection, etc. — using zero-shot and few-shot approaches (i.e., with limited or no custom training data) (Dosovitskiy, 2020; Zohar et al., 2023; Jeeveswaran et al., 2022). In addition, physics-inspired feature engineering has significantly improved model performance by incorporating domain-specific knowledge, leading to more accurate and interpretable predictions in real-world problems (Ompusunggu and Hostens, 2021; Erdmann et al., 2020). In this work, our hypothesis is as follows: Can we leverage state-of-the-art vision foundation models and physics-inspired features, along with transfer learning strategies, to reduce the dependence on site-specific sky camera imagery data?
@@
To address these challenges, we introduce SPIRIT, a novel approach to solar irradiance forecasting with an inductive bias toward enhanced generalizability. In designing, implementing, and evaluating our approach, we make the following contributions:
@@
(1) We develop a novel system that leverages foundation models and physics-informed features, eliminating the need for site-specific model training while enabling effective adaptation across diverse transfer learning scenarios. The flexibility of our framework ensures seamless integration of future advancements in vision models without requiring significant architectural modifications.
@@ 2. Related Work
Traditional methods for solar forecasting have relied heavily on Numerical Weather Prediction (NWP) models and satellite imagery (Markovics and Mayer, 2022). While these methods provide valuable insights, they often lack the spatial and temporal resolution required for accurate short-term forecasts. For instance, NWP models typically operate on a grid scale of several kilometers and update every few hours which may not capture rapid changes in cloud cover that affect solar irradiance (Kostylev et al., 2011). Over the past few years, several time series forecasting approaches have been used for solar forecasting. However, they typically operate on a time frame of multiple hours to day-ahead and are not suitable for capturing short-term variations in solar generation due to transient factors such as cloud cover (Iyengar et al., 2014; Falope et al., 2024).
@@
Use of sky camera imagers for short-term solar forecasting has garnered significant attention in recent years due to their potential to enhance the accuracy of solar power predictions (Hammond et al., 2024; Gao and Liu, 2022; Nie et al., 2024). Sky cameras, equipped with fish-eye lenses, capture wide-angle images of the sky, providing valuable data on cloud cover and movement, which are critical factors in solar irradiance forecasting (Dev et al., 2019). Recent advancements have focused on leveraging sky cameras to address the limitations of traditional approaches. Hammond et al. (Hammond et al., 2024) and Gao et al. (Gao and Liu, 2022) demonstrated the potential of sky cameras in developing high-accuracy models for short-term solar forecasting. These studies utilized extensive site-specific data collected over multiple years to train their models, achieving significant improvements in forecast accuracy compared to traditional methods.
@@
Siddiqui et al.(Siddiqui et al., 2019) proposed a deep learning framework using sky-camera images and auxiliary meteorological data to predict solar irradiance. Their approach employs a convolutional neural network (CNN) with dilated convolutions, followed by an LSTM for temporal forecasting up to four hours ahead. By training on 10 years of data, they demonstrated that incorporating auxiliary data such as temperature, wind speed, and relative humidity enhances generalization and stability in predictions. Similarly, Gao et al. (Gao and Liu, 2022) introduced a transformer-based architecture that integrates a clear sky model to estimate the residual irradiance beyond clear-sky assumptions. Trained on 10 years of data, their model achieves improved forecasting accuracy compared to earlier CNN-LSTM-based methods. Both works underscore the importance of leveraging sky images and auxiliary data for precise solar nowcasting and forecasting. Despite their promise, sky camera-based approaches face challenges related to data availability. With the global solar PV fleet expected to increase from 1 TW in 2022 to 10 TW by 2030, a large majority of the solar farms worldwide will have negligible historical data to train custom models from scratch.
@@
Building upon these challenges, it becomes evident that addressing the limited availability of site-specific data is critical for advancing solar forecasting. Although the use of sky cameras and auxiliary data has substantially improved short-term predictions, the scalability of these methods remains constrained by the dearth of historical data at many solar installations. In this context, transfer learning emerges as a promising solution, as it enables the leveraging of knowledge from pre-trained models and the adaptation of learned representations across different datasets and locations. Notably, previous work such as Nie et al(Nie et al., 2024) has demonstrated that training on a fusion of multiple datasets yields models that perform better on each individual dataset, thereby highlighting the potential benefits of cross-dataset knowledge transfer.
3. SPIRIT Design
3.1. Key Concepts and Problem Setup
@@
Nowcasting refers to the prediction of solar power generation over very short time horizons, typically ranging from a few minutes to a few hours (Lee et al., 2017). In contrast, short-term forecasting extends the prediction horizon to cover periods from one hour to 24 hours (Remund and Müller, 2012). Methods developed to provide forecasts utilize various data sources, such as satellite data (Lopes et al., 2021; Lee et al., 2017), weather station observations (Lee et al., 2017), and sky camera images (Gao and Liu, 2022; Xu et al., 2015; Siddiqui et al., 2019). Nowcasting and short-term forecasting are indispensable for managing the intermittency of solar power, allowing grid operators to perform better scheduling, dispatching, and balancing of energy resources (Dairi et al., 2020; Aouidad and Bouhelal, 2024).
@@
Sky Camera: Sky cameras enhance nowcasting and short-term forecasting by capturing sky images with fish-eye lenses, providing detailed cloud movement, and sun position data. These images enable algorithms to track cloud dynamics and predict their trajectories, essential for estimating solar irradiance (Saraswat et al., 2023; Dev et al., 2019). Offering a low-latency alternative to weather satellites, sky cameras facilitate real-time monitoring. However, variations in camera setup and quality affect image appearance, as shown in Figure 4 in Appendix A.1. As a key tool in solar forecasting, sky cameras contribute to more reliable energy predictions (Rajagukguk et al., 2021). Further details are provided in Appendix B.
Irradiance measurements: Understanding solar irradiance requires distinguishing between three key measurements:
(1) Direct Normal Irradiance (DNI): The amount of solar radiation received per unit area on a surface perpendicular to the sun’s rays without being scattered or diffused by the atmosphere.
(2) Diffuse Horizontal Irradiance (DHI): The portion of solar radiation that reaches a horizontal surface after being scattered by molecules, aerosols, and clouds in the atmosphere. Unlike DNI, DHI comes from all directions in the sky and plays a crucial role during overcast conditions when direct sunlight is obstructed.
(3) Global Horizontal Irradiance (GHI): The total solar radiation received on a horizontal surface, combining both direct and diffuse components. GHI is the sum of DNI, projected onto a horizontal plane, and DHI:
(1) $GHI=DNI\times\cos(\theta)+DHI$
@@
where $\theta$ is the angle between the direction of incoming solar radiation and the vertical, called the zenith angle.
@@
Photovoltaic Power Output: PV power output refers to the electricity generated by solar panels from incoming solar radiation. While it is primarily driven by GHI (Vilanova et al., 2020), factors like temperature, and system losses also play a role. Under stable conditions, the relationship between GHI and PV output is roughly linear (Razak et al., 2016; Natheer Tuaimah and Al-Saidi, 2019). Since PV output is a more actionable metric for grid management and energy planning, predicting it directly is often more desirable.
3.2. Nowcasting Architecture
We propose an architecture that encodes sky images into vector representations, which are augmented with auxiliary data and physics-based features. This representation captures information about the GHI, which is then effectively extracted by a regression model.
Let $\mathcal{X}$ be the set of sky camera images, and $\mathcal{D}$ be the dataset, defined as $\mathcal{D}=\{(X_{i},\mathbf{A}_{i},y_{i})\}_{i=1}^{N}$, where $X_{i}\in\mathcal{X}$ represents the $i$-th sky image, $\mathbf{A}_{i}\in\mathbb{R}^{k}$ corresponds to the auxiliary features such as azimuth and zenith angles of the Sun, and $y_{i}\in\mathbb{R}^{+}$ are the corresponding solar irradiance measurements.
@@ 3.2. Nowcasting Architecture
To leverage domain knowledge in solar power prediction, we introduce a set of additional features, $\mathbf{P}$, derived from the auxiliary measurements $\mathbf{A}$. These features incorporate established solar engineering principles, such as clear sky irradiance, and panel tilt and orientation, as defined in Subsection 3.4. The feature vector $\mathbf{P}$ is given by:
$\mathbf{P}\in\mathbb{R}^{p}$
where $p$ represents the number of physics-based features extracted from the auxiliary data.
@@
where $\oplus$ denotes the concatenation operation. This combined representation leverages data-driven features, visual features, and domain-specific engineering knowledge, providing a comprehensive characterization of each sample $(X_{i},\mathbf{A}_{i},y_{i})\in\mathcal{D}$.
@@
Nowcasting loss function $\mathcal{L}_{nowcast}(\omega)$ is defined as the average of the individual regression losses for each sample, where each individual loss measures the discrepancy between the predicted $\hat{y}_{i}=R_{\omega}(\mathbf{f}_{i})$ and the true value $y_{i}$:
@@
3.3. Forecasting Architecture
Our forecasting architecture processes sequences of sky images to predict GHI across multiple future intervals. Each image is encoded using the embedding and augmentation approach from Section 3.2. A time-series model captures a latent representation of past features, while predictable future covariates, such as the zenith angle, are precomputed and integrated as a vector. The combined past and future representations are then input into a regressor to generate GHI predictions.
A sequence of $T$ images $X_{1:T}=\{X_{1},X_{2},\dots,X_{T}\}$ along with their corresponding auxiliary features $\mathbf{A}_{1:T}=\{\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{T}\}$, where each $\mathbf{A}_{t}\in\mathbb{R}^{k}$ represents the auxiliary feature vector at time $t$, is given.

### v2

# SPIRIT: Short-Term Prediction of Solar IRradIance for Transfer Learning Using Foundation Models

@@ Abstract
Traditional solar forecasting models are based on several years of site-specific historical irradiance data, often spanning five or more years, which are unavailable for newer photovoltaic farms. As renewable energy is highly intermittent, building accurate solar irradiance forecasting systems that are data-efficient is essential for efficient grid management and enabling the ongoing proliferation of solar energy, which is crucial to achieve the United Nations’ net zero goals. In this work, we propose SPIRIT, a novel framework leveraging foundation models for solar irradiance forecasting, making it applicable to newer solar installations. Our approach outperforms state-of-the-art models in zero-shot transfer learning by upto 70%, enabling effective performance at new locations without relying on any past data. Further improvements in performance are achieved through fine-tuning, as more location-specific data becomes available. These findings are supported by statistical significance, further validating our approach. By dramatically reducing the forecasting setup timeline, SPIRIT accelerates solar farm deployment in all potential global sites, most of which lack historical data, thereby democratizing access to clean energy and enabling participation in the renewable energy transition.
Codehttps://github.com/surya-ravindra06/SPIRIT-Official
@@
Figure 1: Illustration of our system: A vision encoder (top-left) extracts embeddings from a sky camera image sampled from a diverse set spanning multiple locations and setups. Physics-inspired features are derived and integrated with auxiliary values, then merged with the image embedding (top-middle) into a unified representation. For nowcasting (right), a regressor predicts Global Horizontal Irradiance from this feature vector. For forecasting (bottom), a time-series model processes past feature vectors to create a context embedding, which is concatenated with a future covariate vector constructed from known future values to form the final latent representation. A regressor then maps this representation to future GHI values (bottom-right).
1 Introduction
@@
The proliferation of solar energy is paramount for electrification and the global energy transition to meet the Net Zero commitments of the United Nations (sadhukhan2022net). Solar energy is notable for its potential to reduce carbon emissions (sen2008solar). Expanding the solar energy infrastructure is crucial to mitigate the effects of climate change (bashir2021enabling) and meet the energy demands arising from sectors such as data centers (agarwal2021redesigning), transportation (lee2016shared), and buildings (iyengar2017cloud).
@@
Unlike conventional sources, solar energy presents unique challenges due to its intermittency from daily and seasonal sunlight variations, significantly impacting grid stability (grid_management). High solar penetration creates the ”duck curve” phenomenon (iyengar2016analyzing), where misalignment between solar production and peak demand complicates grid management. Despite increasing storage capacity, electricity grids function primarily as just-in-time systems requiring balanced energy supply and demand (joskow2012creating). To maintain grid reliability, operators face deviation penalties for unplanned energy contributions (yang2020penalty), necessitating accurate short-term solar predictions for efficient grid operation (iyengar2014solarcast).
@@
Existing approaches for short-term forecasting use sky cameras, a fish-eye lens positioned to look directly towards the zenith, which require extensive site-specific data to train models (hendrikx2024_short_term_forecast; wacv2022). While these methods have demonstrated high accuracy, they rely on multi-year training datasets. With global solar photovoltaic (PV) capacity projected to expand from 1 terawatt in 2022 to 10 terawatts by 2030 (isa2023), 90% of the solar farms worldwide will have negligible data to train custom models from scratch. Thus, lack of site-specific data underscores the need for novel approaches.
@@
Vision foundation models have revolutionized numerous computer vision tasks such as feature extraction and object detection, enabling improved accuracy through zero-shot and few-shot approaches with limited data (google_vit; object_detection_foundation_models; vits_gt_cnns1). Concurrently, physics-inspired feature engineering has enhanced model performance by integrating domain-specific knowledge, yielding more accurate and interpretable predictions (ompusunggu2021physics; erdmann2020physics). Our research hypothesis investigates: Can we leverage state-of-the-art vision foundation models and physics-inspired features, along with transfer learning strategies, to reduce the dependence on site-specific sky camera data?
@@
To address these challenges, we introduce SPIRIT, a novel solar irradiance forecasting approach with inductive bias toward enhanced generalizability. Our contributions include:
@@
(1) Development of an innovative system leveraging foundation models, physics-informed features and future covariates that eliminates site-specific training requirements while facilitating adaptation across diverse transfer learning scenarios. Our system’s flexibility enables seamless integration of future vision model advancements.
@@ 2 Related Work
Traditional solar forecasting methods primarily utilize Numerical Weather Prediction models and satellite imagery (markovics2022comparison). These approaches often lack sufficient spatial and temporal resolution for accurate short-term forecasts. NWP models typically operate at kilometer-scale grids with updates every few hours, failing to capture rapid cloud cover changes affecting solar irradiance (kostylev2011solar). Recent years have seen various time series forecasting approaches for solar prediction, yet these generally function at multi-hour to day-ahead timescales, rendering them inadequate for detecting short-term solar generation variations from transient factors like cloud cover (iyengar2014solarcast).
@@
Sky cameras have gained prominence due to their potential to enhance short-term solar power prediction (wacv2022; yuhao_transfer_learning). These fish-eye lenses capture wide-angle sky images that provide crucial data on cloud cover and movement, key determinants of solar irradiance forecasting (dev2019estimating). Recent research has focused on utilizing sky cameras to overcome prior limitations. wacv2022 and hammond2024 demonstrated sky cameras’ efficacy in developing high-precision short-term solar forecasting models. Their studies employed extensive site-specific data collected across multiple years, yielding models with significantly improved forecast accuracy compared to conventional methods.
@@
talha2019 developed a deep learning framework combining sky-camera imagery with meteorological data for solar irradiance prediction. Their approach integrates a CNN and temporal LSTM for forecasts up to four hours ahead, demonstrating that auxiliary data (temperature, wind speed, relative humidity) enhances prediction generalization when trained on 10 years of data. wacv2022 proposed a transformer-based architecture incorporating a clear sky model to estimate residual irradiance, achieving superior accuracy compared to CNN-LSTM methods using 10 years of training data. Both studies emphasize the efficacy of integrating sky imagery with auxiliary data for precise solar forecasting. However, these approaches face data availability constraints, particularly significant as global solar PV capacity expands from 1 TW (2022) to a projected 10 TW by 2030, leaving most solar installations around the world with insufficient historical data for custom model development.
@@
Limited availability of site-specific data presents a critical challenge for advancing solar forecasting. Despite significant improvements in short-term predictions through sky cameras and auxiliary data integration, these methods face scalability constraints due to insufficient historical data at most solar installations. Transfer learning emerges as a promising solution, enabling knowledge leveraging from pre-trained models and adaptation of learned representations across diverse datasets and locations. Recent research (yuhao_transfer_learning; paletta_transfer_learning) demonstrates that multi-dataset fusion training and data augmentation methods yield models with superior performance, highlighting cross-dataset knowledge transfer benefits in solar forecasting.
3 Design and Implementation
3.1 Key Concepts
@@
Nowcasting and Forecasting: Nowcasting, the prediction of solar power generation over very short time horizons, typically ranges from a few minutes to a few hours (nowcasting_defn). In contrast, short-term forecasting extends the prediction horizon to cover periods from one hour to 24 hours (forecasting_defn). Methods developed to provide forecasts utilize various data sources, such as satellite data (nowcasting_satellite; nowcasting_defn), weather station observations (nowcasting_defn), and sky camera images (wacv2022; talha2019). Nowcasting and short-term forecasting are indispensable for managing the intermittency of solar power, allowing grid operators to perform better scheduling, dispatching, and balancing of energy resources (aouidad2024machine).
@@
Sky Camera: Sky cameras enhance nowcasting and short-term forecasting by capturing sky images with fish-eye lenses, providing detailed cloud movement, and sun position data. These images enable algorithms to track cloud dynamics and predict their trajectories, essential for estimating solar irradiance (saraswat2023sky; dev2019estimating). Offering a low-latency alternative to weather satellites, sky cameras facilitate real-time monitoring. However, variations in camera setup and quality affect image appearance, as shown in Figure 4 in the Appendix. As a key tool in solar forecasting, sky cameras contribute to more reliable energy predictions (rajagukguk2021deep). Further details are provided in Appendix C.
Solar Irradiance necessitates distinguishing between three fundamental measurements. Direct Normal Irradiance (DNI) quantifies solar radiation received per unit area perpendicular to the sun’s rays without atmospheric scattering. Diffuse Horizontal Irradiance (DHI) measures scattered radiation reaching a horizontal surface from all sky directions, becoming crucial during cloud cover. Global Horizontal Irradiance (GHI), the total radiation on a horizontal surface, encompasses both of these components and is mathematically expressed as:
$GHI=DNI\times\cos(\theta)+DHI$ (1)
@@
where $\theta$ represents the zenith angle, defined as the angle between the direction of incoming solar radiation and the vertical.
@@
Photovoltaic Power Output refers to the electricity generated by solar panels from incoming solar radiation. While it is primarily driven by GHI (ghi_pv_linear), factors like temperature, and system losses also play a role. Under stable conditions, the relationship between GHI and PV output is roughly linear (pv_output). Since PV output is a more actionable metric for grid management and energy planning, predicting it directly is often more desirable.
3.2 Physics-inspired Feature Engineering
Clear sky models (ineichen1; clearsky1) are mathematical models that estimate the solar irradiance at a given location under cloud-free conditions, serving as a representation of the maximum possible radiation reaching the Earth’s surface. These models leverage fundamental atmospheric physics and employ mathematical formulations based on solar geometry, atmospheric transmittance, and radiative transfer (clearsky1). The Ineichen model (ineichen1) requires inputs such as latitude, longitude, time, and date, which are readily available. This allows clear sky irradiance values to be readily computed and incorporated into our model as features, providing a reference for expected irradiance levels in the absence of cloud interference.
Physics behind solar irradiance: Solar irradiance varies significantly based on site-specific parameters like panel orientation and solar positioning. The critical angle of incidence ($\theta$), representing the angle between incoming solar rays and panel normal, is calculated as $\cos(\theta)=\cos(\theta_{z})\cdot\cos(\beta)+\sin(\theta_{z})\cdot\sin(\beta)\cdot\cos(\gamma-\alpha)$, where $\theta_{z}$ and $\gamma$ denote solar zenith and azimuth angles, while $\beta$ and $\alpha$ represent panel tilt and azimuth angles, respectively.
These physics-inspired features enable the creation of future covariates through their deterministic calculation capabilities independent of temporal constraints, thereby enhancing time-series information for predictive modeling.
3.3 Nowcasting Architecture
We propose an architecture that encodes sky images into vector representations, which are augmented with auxiliary data and physics-based features to capture GHI information that is then effectively extracted by a regression model. Let $\mathcal{X}$ be the set of sky camera images, and $\mathcal{D}$ be the dataset, defined as $\mathcal{D}=\{(X_{i},\mathbf{A}_{i},y_{i})\}_{i=1}^{N}$, where $X_{i}\in\mathcal{X}$ represents the $i$-th sky image, $\mathbf{A}_{i}\in\mathbb{R}^{k}$ corresponds to the auxiliary features such as azimuth and zenith angles of the Sun, and $y_{i}\in\mathbb{R}^{+}$ are the corresponding solar irradiance measurements.
@@ 3.3 Nowcasting Architecture
To leverage domain knowledge in solar power prediction, we introduce a set of additional features, $\mathbf{P}$, derived from the auxiliary measurements $\mathbf{A}$. These features incorporate established solar engineering principles, such as clear sky irradiance, and panel tilt and orientation, as defined in Subsection 3.2. The feature vector is given by $\mathbf{P}\in\mathbb{R}^{p}$ where $p$ represents the number of physics-based features extracted from the auxiliary data.
@@
where $\oplus$ denotes concatenation. This combined representation leverages data-driven, visual and domain-specific features, providing a comprehensive characterization of each sample $(X_{i},\mathbf{A}_{i},y_{i})\in\mathcal{D}$.
@@
Nowcasting loss function $\mathcal{L}_{nowcast}(\omega)$ is defined as the average of the individual regression losses for each sample, where each individual loss measures the discrepancy between the predicted $\hat{y}_{i}=R_{\omega}(\mathbf{f}_{i})$ and the true value $y_{i}$ across all $N$ samples in the dataset:
@@
3.4 Forecasting Architecture
Our forecasting architecture processes sequences of sky images to predict GHI across multiple future intervals. Each image is encoded using the embedding and augmentation approach from Section 3.3, creating a foundation for temporal analysis. A sequence of $T$ images $X_{1:T}=\{X_{1},X_{2},\dots,X_{T}\}$ along with their corresponding auxiliary features $\mathbf{A}_{1:T}=\{\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{T}\}$, where each $\mathbf{A}_{t}\in\mathbb{R}^{k}$ represents the auxiliary feature vector at time $t$, is given.